OpenStudy (anonymous):

What is the solution of Log_2x+3^125=3?

Nnesha (nnesha):

\[\huge\rm log_{2x+3}125=3\] like this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

log(2x-3) 125 = 3 (2x - 3)^3 = 125 ........log(baseA)B = x , then A^x = B Taking cube roots on both sides 2x - 3 = 5 2x = 5+3 2x = 8 x = 4

Nnesha (nnesha):

alright so you have to convert log to exponential form here is an example